Simon Company’s year-end balance sheets follow.

At December 31 2017 2016 2015

Assets Cash $ 32,000 $ 34,900 $ 37,200

Accounts receivable, net 89,600 62,600 58,400

Merchandise inventory 66,406 83,500 50,900

Prepaid expenses 11,305 10,389 4,376

Plant assets, net 395,689 293,611 239,124

Total assets $ 595,000 $ 485,000 $ 390,000

Liabilities and Equity Accounts payable $ 149,637 $ 81,965 $ 51,480

Long-term notes payable secured by mortgages on plant assets 112,979 112,666 84,466

Common stock, $10 par value 162,500 162,500 162,500

Retained earnings 169,884 127,869 91,554

Total liabilities and equity $ 595,000 $ 485,000 $ 390,000

The company’s income statements for the years ended December 31, 2017 and 2016, follow.

For Year Ended December 31 2017 2016

Sales $ 773,500 $ 577,150

Cost of goods sold $ 471,835 $ 375,148

Other operating expenses 239,785 146,019

Interest expense 13,150 13,274

Income taxes 10,056 8,657

Total costs and expenses 734,826 543,098

Net income $ 38,674 $ 34,052

Earnings per share $ 2.38 $ 2.10

Additional information about the company follows. Common stock market price, December 31, 2017 $ 31.00 Common stock market price, December 31, 2016 29.00 Annual cash dividends per share in 2017 0.36 Annual cash dividends per share in 2016 0.18 To help evaluate the company's profitability, compute the following ratios for 2017 and 2016:

1. Return on common stockholders' equity.

2. Price-earnings ratio on December 31.

3. Dividend yield.
Answer #1

Calculate following

2017 2016
Return on Common Stockholder's equity Net income/Average common equity 38674/311376.50 = 12.42% 34052/272211.50 = 12.51%
Price earning ratio MPS/EPS 31/2.38 = 13.03 29/2.10 = 13.81
Dividend yield DPS/MPS 0.36/2.38 = 15.13% 0.18/2.10 = 8.57%